A Geiger counter counts the number of alpha particles from radioactive material. Over a long period of time, an average of 5 particles per minute occurs. Assume the arrival of particles at the counter follows a Poisson distribution. Round your answers to 3 decimals.

a. Find the probability that no particle arrives in a particular one minute period.

b. Find the probability that at least one particle arrive in a particular one minute period.
Answer #1

This is a Poisson distribution with parameter:

\lambda = 5

Hence,

a) P(X = 0)

= \frac{e^{-5}5^0}{0!}

= 0.007

b) P(Atleast one)

= 1 - P(None)

= 1 - 0.007

= 0.993